What is answer?
A group of boy scout can row 4km in still water per hour. it takes them as long to 6km upstream as 10 km downstream. How fast it is current?

Answer by mananth(12270)

Boat speed = 4 kph
current speed = x kph

speed against current 4 - x
speed with current 4 + x

Distance against current 6 km
Distance with current 10 km

t=d/r time against = time with
6/(4-x)=10/(4+x )
6(4+x)= 10(4-x )
24x=40-10x
10x+6x = 16
16 x = 16
/ 16
x = 1 kph current speed
m.ananth@hotmail.ca